OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line that passes through (4, 2) with a slope of 3.

OpenStudy (anonymous):

Slope intercept form of the line is y = mx + b Substituting the given values of y, m and x (2) = (3)(4) + b 2 = 12 + b 2 - 12 = b - 10 = b Then the line equation must be y = 3x – 10

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

can you answer this?????????? http://openstudy.com/study#/updates/4f1dd736e4b04992dd24586f

OpenStudy (anonymous):

u r welcome....☺ let me check for the second one....